OpenStudy (anonymous):

Solve by using the quadratic formula to solve x2 + 6x = -3.

OpenStudy (mathmale):

jonny, hope you have the "quadratic formula" written down for reference somewhere. If not, please write it down now and keep this note handy:\[IF~ax^2+bx+c=0,~then~x=\frac{ -b \pm \sqrt{b^2-4ac}}{ 2a }\]

OpenStudy (mathmale):

Here ax^2+bx+c=0 is a quadratic equation, and its solutions can be found through the 2nd formula, the quadratic formula.

OpenStudy (mathmale):

Please write x2 + 6x = -3 as\[x^2+6x+3=0.\]

OpenStudy (mathmale):

Can you explain why I moved the 3? Can you explain why I changed your x2 to \[x^2?\]

OpenStudy (anonymous):

you moved the 3 to make the equation equal to 0 im not sure why you changed the x2

OpenStudy (mathmale):

Right. Your equation MUST be in the form ax^2+bx+c=0, so that you can easily identify the values of a, b and c.

OpenStudy (mathmale):

x2 could be ambiguous (confusing); some people would interpret it as \[x _{2},\]whereas others might interpret it as I did, as \[x^2. \]

OpenStudy (mathmale):

If you don't want to use equation Editor, you could write "x squared" as x^2.

OpenStudy (mathmale):

jonny, look at your \[x^2+6x+3=0\]and compare it to \[ax^2+bx+c=0.\]

OpenStudy (mathmale):

Can you now identify the values of a, b and c?

OpenStudy (mathmale):

a=? b=? c=? 0=0

OpenStudy (anonymous):

a=x b=6 c=3

OpenStudy (anonymous):

a=1

OpenStudy (mathmale):

Cool! b=?

OpenStudy (anonymous):

b=6

OpenStudy (mathmale):

Right on! c=?

OpenStudy (anonymous):

c=3

OpenStudy (mathmale):

Perfect. So, now, jonny, you need to copy down that quadratic formula\[x=\frac{ -b \pm \sqrt{b^2-4ac}}{ 2a }\]and substitute your values for a, b and c. Bet you've done this before.

OpenStudy (anonymous):